# An Effective Way to Determine the Separability of Quantum State

### v1

@@ Abstract
We propose in this work a practical approach, by virtue of correlation matrices of the generic observables, to study the long lasting tough issue of quantum separability. Some general separability conditions are set up through constructing a measurement-induced Bloch space. In essence, these conditions are established due to the self constraint in the space of quantum states. The new approach can not only reproduce many of the prevailing entanglement criteria, but also lead to even stronger results and manifest superiority for some bound entangled states. Moreover, as a by product, the new criteria are found directly transformable to the entanglement witness operators.
Introduction
@@
Quantum entanglement captures the most enigmatic and deepest insights into quantum theory, and has deep connections with other branches of physics [1]. In quantum information theory, how to distinguish an entangled state from a separable state is a fundamental and significant problem, say the so-called separability problem [2, 3]. It is well-known that the separability problem tends to be a NP-hard issue with the increase of dimension, even for the bipartite system [4], to develop a practical and experimental friendly separability criterion is pivotal in the establishment of a complete entanglement theory [2, 3]. In past decades, numerous methods have been put forward to attack it, including positive map theory and entanglement witness [5, 6, 7], the computable cross norm or realignment (CCNR) criterion [8, 4, 3], correlation matrix method [11, 12, 13, 5], local uncertainty relation (LUR) criteria [15, 16, 17, 18, 19, 20], covariance matrix [21, 22], quantum Fisher information (QFI) method [23, 24, 25], moment method [26, 27, 28, 29], etc.
In this Letter, we develop a framework to derive separability criterion for finite-dimensional systems in terms of the correlation matrix. To realize it, a measurement-induced, generalized Bloch space for the arbitrary observables is introduced, and by which a generic separability criterion is formulated. The key point here is that the correlation matrix of the separable state is limited by the size of the generalized Bloch space. The powerfulness of this approach can be realized by noticing that in which the de Vicente’s correlation matrix criterion [11], CCNR criterion [4, 3], the correlation matrix criterion based on symmetric informationally complete positive operator-valued measures (SIC-POVM) [5] etc. can all be reproduced transparently, and even stronger criterion may be obtained. Moreover, the new scenario also provides a concise method to construct an entanglement witness operator via generic observables. Examples are provided to show the availability of bound entangled state detection.
Bloch representation of observables and density matrix
@@
The notations to be employed in the discussion go as follows. The observable is denoted by a capital letter, such as $A$, and a boldface capital letter $\boldsymbol{A}$ implies a measurement vector containing certain observables, i.e. $\boldsymbol{A}=(A_{1},A_{2},\cdots)$. The density matrix $\rho$ of a $d$-level quantum system is formulated in Bloch representation, to wit [30, 31, 32, 33, 35, 34]
@@
where $t_{\mu}=\mathrm{Tr}[A_{\mu}]$ and $a_{\mu\nu}=(\vec{a}_{\mu})_{\nu}=\mathrm{Tr}[A_{\mu}\pi_{\nu}]/2$. And then, a measurement vector $\boldsymbol{A}$ can be parameterized as a real matrix $M$ in Bloch representation
$\boldsymbol{A}\rightarrow M_{A}\equiv\left(\begin{matrix}\frac{t_{1}}{\sqrt{2d}}&\frac{t_{2}}{\sqrt{2d}}&\cdots\\ \vec{a}_{1}&\vec{a}_{2}&\cdots\end{matrix}\right)\;.$ (5)
@@
Correlation matrix of separable state and measurement-induced Bloch space
Quantum entanglement exhibits inherent non-classical correlation in many-body systems, even far apart from each other, which was initially recognized by Einstein, Podolsky and Rosen [36] and Schrödinger [37]. Based on correlation function [38] Bell inequality provides a practical way for the first time to test the entanglement. In the family of Bell inequality, the Clauser-Horne-Shimony-Holt (CHSH) inequality [39]
$\left|\braket{A\otimes B}-\braket{A\otimes B^{\prime}}+\braket{A^{\prime}\otimes B}+\braket{A^{\prime}\otimes B^{\prime}}\right|\leq 2\;$ (6)
might be the most prominent one, which tends to be experiment-operational friendly. Here, $A(A^{\prime})$ and $B(B^{\prime})$ are pairs of dichotomic observables with eigenvalues of ±$\pm 1$ measured by the parties $A$ and $B$, respectively. The violation of CHSH inequality and the predictions of quantum theory were first confirmed experimentally by [40, 41, 42], then later more convincingly by some loophole-free tests [43, 44, 45, 46]. The violation of CHSH inequality implies the entanglement and exclusion of local hidden variable scenario for the description of the correlation system. However, in reverse, Bell inequalities fail to recognize all the entangled states [47]. Hence, to explore and exploit the nature of entanglement, it is essential to find even stronger criteria for the determination of separable state.
To this end, we define the following correlation matrix entry for arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$
@@
Generally speaking, a state is separable iff it is a convex combination of product states, i.e. $\rho_{\mathrm{sep}}=\sum_{i}q_{i}\rho_{i}^{A}\otimes\rho_{i}^{B}$, where $\{\rho_{i}^{A}\}$ and $\{\rho_{i}^{B}\}$ are two pure subsystems. And, hence the corresponding correlation matrix for a separable state can write
@@
where $\braket{A_{\mu}}_{i}=\mathrm{Tr}[\rho_{i}^{A}A_{\mu}]$ and $\braket{B_{\nu}}_{i}=\mathrm{Tr}[\rho_{i}^{B}B_{\nu}]$. Eq. 12 inspires us to define a measurement-induced Bloch vector(MIBV) $\vec{\alpha}$ for measurements $\boldsymbol{A}$, of which
@@
with $\bar{A}_{\mu}=\vec{a}_{\mu}\cdot\vec{\pi}$ denoting the traceless part of observable $A_{\mu}$. And, various MIBV $\vec{\alpha}$ span the corresponding measurement-induced Bloch space, signified as $\mathcal{B}(\boldsymbol{A})$, which has been applied to the study of certainty and uncertainty relations [1]. The correlation matrix of a separable state can be reformulated by MIBVs, like
@@
Here, $\vec{t}^{A}$ and $\vec{t}^{B}$ are column vectors consisting of traces of measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$, respectively; $\vec{\alpha}_{i}$ and $\vec{\beta}_{i}$ are MIBVs of measurements $\boldsymbol{A}$ and $\boldsymbol{B}$; $\mathrm{T}$ denotes the transpose of a given matrix.
The separability criteria
A family of sufficient conditions for entanglement of measurements $\boldsymbol{A}$ and $\boldsymbol{B}$ goes as follows, with proof given in the Supplement:
@@
Let $\mathcal{C}$ be a correlation matrix for arbitrary measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, if the corresponding state $\rho$ is separable, then
@@
Otherwise, $\rho$ will be an entangled state. Here, $\|X\|_{\operatorname{tr}}=\mathrm{Tr}\left[\sqrt{X^{\dagger}X}\right]$ is the trace norm of matrix $X$; $\kappa_{A}=\sqrt{(\vec{t}\,^{A})^{2}/d_{A}^{2}+\frac{d_{A}-1}{d_{A}}\lambda_{\max}(\Omega_{A})}$ with $\lambda_{\max}(X)$ being the maximal eigenvalue of matrix $X$ and $(\Omega_{A})_{\mu\nu}=2\vec{a}_{\mu}\cdot\vec{a}_{\nu}$, and similarly for $\kappa_{B}$.
Next, we illustrate how Theorem 1 works. For a bipartite system, we have the following Bloch representation
@@
Here, $\Pi_{0}^{X}=\sqrt{\frac{2}{d_{X}}}\mathds{1}$; $\{\Pi_{\mu}^{X}=\pi_{\mu}^{X}\}_{\mu=1}^{d^{2}-1}$ are generators of $\mathfrak{su}(d_{X})$ Lie algebra for $X=A,\ B$; the coefficient matrix $\chi_{\mu\nu}=\mathrm{Tr}[\rho(\Pi_{\mu}^{A}\otimes\Pi_{\nu}^{B})]$; $a_{\mu}=\mathrm{Tr}[\rho\pi_{\mu}^{A}\otimes\mathds{1}]$; $b_{\mu}=\mathrm{Tr}[\rho\mathds{1}\otimes\pi_{\mu}^{B}]$; $\mathcal{T}_{\mu\nu}=\mathrm{Tr}[\rho\pi_{\mu}^{A}\otimes\pi_{\nu}^{B}]$; and $\chi_{00}=\frac{2}{\sqrt{d_{A}d_{B}}}$ since $\mathrm{Tr}[\rho]=1$. The correlation matrix can then be reexpressed as
@@
with $M_{A}$ and $M_{B}$ being the Bloch representations of measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$ defined in Eq. 5. Equipped with this, one may notice that the traceless part of the measurement $\boldsymbol{A}$ can be viewed as the vertex of a polygon or polyhedron in $(d^{2}-1)$-dimensional space, which is enlightening for further processing of Theorem 1.
Consider the simplest situation $t_{\mu}^{A}=0$, and $\vec{a}_{1}=\vec{0}$, $\vec{a}_{\mu}=(0,\cdots,0,1,0,\cdots,0)^{\mathrm{T}}$, $\mu=2,\cdots,d_{A}^{2}$, where all components of $\vec{a}_{\mu}$ equal to 00, except the $\mu$th equals to $1$. It is easy to find $\lambda_{\max}(\Omega_{A})=2$ (similarly for $B$) and the separability condition
@@
Here, $\mathcal{T}=\mathrm{Tr}[\rho(\pi_{\mu}\otimes\pi_{\nu})]$. Note, Eq.(19) simply reproduces the de Vicente’s correlation matrix criterion [11].
Moreover, if we let $t_{1}^{A}=\sqrt{2d}h_{A},t_{1}^{B}=\sqrt{2d}h_{A},t_{\mu}^{A}=t_{\mu}^{B}=0,\mu\neq 1$, and take $\{\vec{a}_{\mu}\}$ and $\{\vec{b}_{\mu}\}$ the same as in above, we obtain
$\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{\frac{2(d_{A}-1+h_{A}^{2})}{d_{B}}}\sqrt{\frac{2(d_{B}-1+h_{B}^{2})}{d_{B}}}\;,$ (20)
which is exactly the Sarbicki et al.’s criterion [6]. Whereas, if $\boldsymbol{A}$, and similarly for $\boldsymbol{B}$, has the same trace and its traceless part corresponds to a regular simplex, viz.
$M_{A}=\left(\begin{matrix}\frac{t_{A}}{\sqrt{2d_{A}}}&\frac{t_{A}}{\sqrt{2d_{A}}}&\cdots&\frac{t_{A}}{\sqrt{2d_{A}}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right)\;,$ (23)
where $\{\vec{a}_{\mu}\}$ constitutes a regular $(d_{A}^{2}-1)$-simplex, i.e. $\left\{\vec{a}_{\mu}\Big{|}\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}\right\}$ for $\mu,\nu=1,\cdots,d_{A}^{2}$, a family of separability criteria can be readily obtained (the proof of separability criteria and construction of regular simplex vertex coordinates are given in Supplements):
@@
Given measurement vectors $\boldsymbol{A}$ and $\boldsymbol{B}$, let their trace parts satisfy $t_{\mu}^{A}=t_{A},t_{\mu}^{B}=t_{B}$ and traceless parts correspond respectively to a regular simplex. The correlation matrix of separable state $\rho$ then satisfy
@@
The 1 represents a family of separability conditions for different $t_{A}$ and $t_{B}$, many of existing criteria appear to be its particular cases. When $t_{X}$ takes $\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ and $\sqrt{\frac{2d_{X}}{d_{X}-1}}$ ($X=A,B$), it gives out the well-known CCNR [50, 4] and entanglement criterion via SIC-POVM (ESIC) criteria [5], respectively. To exhibit more, following we compare 1 with some symbolic criteria under certain quantum states, the Horodecki’s bound entangled states with white noise $\rho_{\mathrm{H}}(s,p)$, unextendible product bases (UPB) bound entangled states with white noise $\rho_{\mathrm{UPB}}(p)$ and chessboard states $\rho_{\mathrm{CB}}$. The exact forms of these quantum states are shown in the Supplemental Material.
In Fig. 1 we show and compare 1, CCNR and ESIC criteria under three typical classes of bound entangled states, where larger $t$ witnesses higher robustness against white noise and hence at which larger portion of entangled state may be detected. Fig. 1(a) exhibits Horodecki’s bound entangled states with white noise and Fig. 1(b) explains the detection results for UPB state with white noise (peru solid line) and $50000$ randomly generated chessboard state (blue dashed line). Note, as shown in the figure, 1 reproduces the CCNR and ESIC criteria as two special cases when $t=\sqrt{3}/2$ and $t=\sqrt{3}$, respectively.

### v2

@@ Abstract
We propose in this work a practical approach to address the longstanding and challenging problem of quantum separability, leveraging the correlation matrices of generic observables. General separability conditions are obtained by dint of constructing the measurement-induced Bloch space, which in essence come from the intrinsic constraints in the space of quantum state. The novel approach can not only reproduce various established entanglement criteria, it may as well brings about some new results, possessing obvious advantages for certain bound entangled states and the high dimensional Werner states. Moreover, it is found that criteria obtained in our approach can be directly transformed into entanglement witness operators.
1 Introduction
@@
Quantum entanglement captures the most enigmatic and deepest insights of quantum theory, and has a broad connections with other branches of physics [1]. In quantum information theory, distinguishing entangled states from separable states appears to be a fundamental and challenging issue, commonly referred to as the separability problem [2, 3]. It is widely recognized that the separability problem becomes NP-hard as the dimension increases, even in bipartite systems [4]. Therefore, developing a practical and experimentally feasible separability criterion is crucial for establishing a comprehensive entanglement theory [2, 3]. In past decades, a multitude of approaches have been proposed to address this issue, including positive map theory and entanglement witnesses [5, 6, 7], spin squeezing inequalities [8, 9, 10, 11], the computable cross norm or realignment (CCNR) criterion [12, 13, 14], correlation matrix method [15, 16, 17, 18, 19], local uncertainty relation (LUR) criteria [20, 21, 22, 23, 24, 25], covariance matrix [26, 27], quantum Fisher information (QFI) method [28, 29, 30, 31], moment method [32, 33, 34, 35], etc.
In this paper, we propose a comprehensive framework to establish a separability criterion for finite-dimensional quantum systems using the correlation matrix. To achieve this, we introduce a measurement-induced generalized Bloch space for arbitrary observables, which enables us to formulate a generic separability criterion. A critical observation is that the correlation matrix of separable states is constrained by the scale parameter of the generalized Bloch space. The versatility and strength of this approach are evident from its ability to reproduce several established criteria, such as de Vicente’s correlation matrix criterion [15], the CCNR criterion [13, 14], the correlation matrix criterion based on symmetric informationally complete positive operator-valued measures (SIC-POVM) [18] and Sarbicki et al.’s criterion [19], while the potential to yield some other even stronger results. Additionally, this framework provides a streamlined method for constructing entanglement witness operators using generic observables. We provide examples demonstrating the effectiveness of detecting bound entangled states and the high-dimensional Werner states.
2 Bloch representation of observables and density matrix
@@
Before proceeding, we introduce the necessary notations. An observable is denoted by a capital letter, such as $A$, while a boldface capital letter $\boldsymbol{A}$ represents an $m$-tuple of observables, i.e., $\boldsymbol{A}=(A_{1},A_{2},\cdots)$. The density matrix $\rho$ of a $d$-level quantum system is expressed in the Bloch representation, specifically [36, 37, 38, 39, 40, 41]
@@
where $t_{\mu}=\mathrm{Tr}[A_{\mu}]$ and $a_{\mu\nu}=(\vec{a}_{\mu})_{\nu}=\mathrm{Tr}[A_{\mu}\pi_{\nu}]/2$. And then, an $m$-tuple $\boldsymbol{A}$ of observables can be parameterized as a real matrix $M$ in Bloch representation
$\boldsymbol{A}\rightarrow M_{A}\equiv\left[\begin{matrix}\frac{t_{1}}{\sqrt{2d}}&\frac{t_{2}}{\sqrt{2d}}&\cdots\\ \vec{a}_{1}&\vec{a}_{2}&\cdots\end{matrix}\right]\;.$ (3)
@@
3 Correlation matrix of separable state and measurement-induced Bloch space
Quantum entanglement exhibits inherent non-classical correlations in many-body systems, even when the particles are far apart. This phenomenon was initially recognized by Einstein, Podolsky, and Rosen [42], as well as by Schrödinger [43]. Based on correlation functions, Bell inequalities provide a practical method to test entanglement for the first time [44]. Among the family of Bell inequalities, the Clauser-Horne-Shimony-Holt (CHSH) inequality [45]
$\left|\!\braket{A\otimes B}\!\!-\!\!\braket{A\otimes B^{\prime}}\!\!+\!\!\braket{A^{\prime}\otimes B}\!\!+\!\!\braket{A^{\prime}\otimes B^{\prime}}\!\right|\!\leq\!2$ (4)
is perhaps the most prominent. Here, $A(A^{\prime})$ and $B(B^{\prime})$ are pairs of dichotomic observables with eigenvalues of ±$\pm 1$ measured by the parties $A$ and $B$, respectively. The violation of CHSH inequality and the predictions of quantum theory were first confirmed experimentally by [46, 47, 48], then later more convincingly by some loophole-free tests [49, 50, 51, 52]. Although the violation of the CHSH inequality certifies entanglement and excludes the possibility of a local hidden variable description for correlated systems, Bell inequalities inherently fail to detect all entangled states [53]. Therefore, to fully explore and utilize the nature of entanglement, it is essential to establish even stronger criteria for determining separable states.
To this end, we define the following correlation matrix entries for arbitrary $m$-tuples $\boldsymbol{A}$ and $\boldsymbol{B}$
@@ 3 Correlation matrix of separable state and measurement-induced Bloch space
A state is separable iff it is a convex combination of product states, i.e. $\rho_{\mathrm{sep}}=\sum_{i}q_{i}\rho_{i}^{A}\otimes\rho_{i}^{B}$, where $\{\rho_{i}^{A}\}$ and $\{\rho_{i}^{B}\}$ are two pure subsystems [53]. Hence the correlation matrix of a separable state can be written as
@@
where $\braket{A_{\mu}}_{i}=\mathrm{Tr}[\rho_{i}^{A}A_{\mu}]$ and $\braket{B_{\nu}}_{i}=\mathrm{Tr}[\rho_{i}^{B}B_{\nu}]$. Eq. 6 inspires us to define a measurement-induced Bloch vector (MIBV) $\vec{\alpha}$ for $\boldsymbol{A}$, given by
@@
where $\bar{A}_{\mu}=\vec{a}_{\mu}\cdot\vec{\pi}$ denotes the traceless part of the observable $A_{\mu}$. Various MIBV $\vec{\alpha}$ span the corresponding measurement-induced Bloch space (MIBS), denoted as $\mathcal{B}(\boldsymbol{A})$, which is defined as
$\mathcal{B}(\!\boldsymbol{A}\!)\!=\!\{\vec{\alpha}\!=\!(\braket{\!\bar{A}_{1}\!}_{\rho},\!\cdots,\!\braket{\!\bar{A}_{m}\!}_{\rho})^{\mathrm{T}}\!|\rho\in\!\mathcal{S}(\mathcal{H})\!\}\;.$ (8)
Here, $\mathcal{S}(\mathcal{H})$ is the set of density matrices. It is noteworthy that $\mathcal{B}(\boldsymbol{A})$ is also referred to as the joint numerical range (JNR) of observables $\bar{A}_{1},\cdots,\bar{A}_{m}$ [54, 55], which has been extensively studied in the context of uncertainty relations, quantum entanglement, and more [25, 56]. In particular, the quantity
$\beta\equiv\sup_{\rho}\sum_{k=1}^{m}\braket{\bar{A}_{k}}_{\rho}^{2}=\max_{\vec{x}\in\mathcal{B}(\boldsymbol{A})}|\vec{x}|^{2}$ (9)
is a critical parameter of JNR or MIBS, closely related to state-independent uncertainty relation [57, 58]. In the following sections, we will establish an explicit relationship between $\beta$ and the quantum separability problem. The correlation matrix of a separable state can be reformulated using MIBVs, as follows:
@@
Here, $\vec{t}^{A}$ and $\vec{t}^{B}$ are column vectors consisting of the traces of measurements $\boldsymbol{A}$ and $\boldsymbol{B}$, respectively; $\vec{\alpha}_{i}$ and $\vec{\beta}_{i}$ are MIBVs of measurements $\boldsymbol{A}$ and $\boldsymbol{B}$; and $\mathrm{T}$ denotes the transpose of a given matrix.
4 The separability criteria
With this, we are now ready to present the main results of this work, with proof given in the Appendix A:
@@ 4 The separability criteria
Let $\mathcal{C}$ be a correlation matrix for arbitrary $\boldsymbol{A}$ and $\boldsymbol{B}$ with $\sum_{\mu}t_{\mu}^{A}\vec{a}_{\mu}=\sum_{\mu}t_{\mu}^{B}\vec{b}_{\mu}=0$, if the corresponding state $\rho$ is separable, then
@@
Otherwise, $\rho$ will be an entangled state. Here, $\|X\|_{\operatorname{tr}}=\mathrm{Tr}\left[\sqrt{X^{\dagger}X}\right]$ is the trace norm of matrix $X$; $\kappa_{A}=\sqrt{(\vec{t}^{\,A})^{2}/d_{A}^{2}+\beta_{A}}$ with $\beta_{A}=\sup_{\rho}\sum_{k=1}^{m}\braket{A_{k}}_{\rho}^{2}$, and similarly for $\kappa_{B}$.
Note, Theorem 1 provides an alternative approach to determining the separability of quantum states, different from the covariance matrix (CM) criteria [26, 27]. The CM criteria make use of the concavity and the positive definitiveness of CM, while Theorem 1 stems from the geometry property of the MIBS, highlighting the independence of the covariance matrix and correlation matrix methods. Next, we illustrate how Theorem 1 works. For a bipartite system, we have the following Bloch representation
@@
Here, $\Pi_{0}^{X}=\sqrt{\frac{2}{d_{X}}}\mathds{1}$; $\{\Pi_{\mu}^{X}=\pi_{\mu}^{X}\}_{\mu=1}^{d^{2}-1}$ are the generators of the $\mathfrak{su}(d_{X})$ Lie algebra for $X=A,B$; coefficient matrix is given by $\chi_{\mu\nu}=\mathrm{Tr}[\rho(\Pi_{\mu}^{A}\otimes\Pi_{\nu}^{B})]$; $a_{\mu}=\mathrm{Tr}[\rho\pi_{\mu}^{A}\otimes\mathds{1}]$; $b_{\mu}=\mathrm{Tr}[\rho\mathds{1}\otimes\pi_{\mu}^{B}]$; $\mathcal{T}_{\mu\nu}=\mathrm{Tr}[\rho\pi_{\mu}^{A}\otimes\pi_{\nu}^{B}]$; and $\chi_{00}=\frac{2}{\sqrt{d_{A}d_{B}}}$ since $\mathrm{Tr}[\rho]=1$. The correlation matrix can then be expressed as
@@
where $M_{A}$ and $M_{B}$ are the Bloch representations of $\boldsymbol{A}$ and $\boldsymbol{B}$ defined in Eq. 3. With this framework, one may observe that the traceless part of the measurement $\boldsymbol{A}$ can be interpreted as the vertices of a polygon or polyhedron in $(d^{2}-1)$-dimensional space, which provides valuable insights for the further analysis of Theorem 1.
Let us consider the simplest scenario where $t_{\mu}^{A}=0$, $\vec{a}_{1}=\vec{0}$, and $\vec{a}_{\mu}=(0,\cdots,0,1,0,\cdots,0)^{\mathrm{T}}$ for $\mu=2,\cdots,d_{A}^{2}$. Here, all components of $\vec{a}_{\mu}$ are zero except for the $(\mu-1)$th component which equals 1. It is straightforward to determine that $\beta_{A}=\frac{2(d_{A}-1)}{d_{A}}$ (analogously for $\beta_{B}$). The separability condition can be expressed as:
@@
Notably, Eq.(15) simply reproduces the de Vicente’s correlation matrix criterion [15].
Furthermore, by setting $t_{1}^{A}=\sqrt{2d_{A}}h_{A}$, $t_{1}^{B}=\sqrt{2d_{B}}h_{B}$, and $t_{\mu}^{A}=t_{\mu}^{B}=0$ for $\mu\neq 1$, and choosing $\{\vec{a}_{\mu}\}$ and $\{\vec{b}_{\mu}\}$ as previously defined, we obtain
$\|\mathcal{C}\|_{\mathrm{tr}}\leq\sqrt{\frac{2(d_{A}-1+h_{A}^{2})}{d_{A}}}\sqrt{\frac{2(d_{B}-1+h_{B}^{2})}{d_{B}}},$
which is precisely the Sarbicki et al.’s criterion [19]. Additionally, if $\boldsymbol{A}$ (and similarly for $\boldsymbol{B}$) has the same trace and its traceless part corresponds to a regular simplex, specifically,
$M_{A}=\left[\begin{matrix}\frac{t_{A}}{\sqrt{2d_{A}}}&\frac{t_{A}}{\sqrt{2d_{A}}}&\cdots&\frac{t_{A}}{\sqrt{2d_{A}}}\\ \vec{a}_{1}&\vec{a}_{2}&\cdots&\vec{a}_{d_{A}^{2}}\end{matrix}\right]\;,$ (16)
where $\{\vec{a}_{\mu}\}$ forms a regular $(d_{A}^{2}-1)$-simplex, i.e., $\vec{a}_{\mu}\cdot\vec{a}_{\nu}=\frac{d_{A}^{2}\delta_{\mu\nu}-1}{d_{A}^{2}-1}$ for $\mu,\nu=1,\ldots,d_{A}^{2}$, a family of separability criteria can be readily derived. The proof and the construction of the regular simplex vertex coordinates are provided in Appendices B and C.
@@
Given $\boldsymbol{A}$ and $\boldsymbol{B}$, let their trace parts satisfy $t_{\mu}^{A}=t_{A},t_{\mu}^{B}=t_{B}$ and traceless parts correspond respectively to a regular simplex. The correlation matrix of separable state $\rho$ then satisfies
@@
The 1 represents a family of separability conditions for various values of $t_{A}$ and $t_{B}$, with many existing criteria appearing as particular cases of this family. Specifically, when $t_{X}$ takes the values $\sqrt{\frac{2d_{X}}{d_{X}^{2}-1}}$ and $\sqrt{\frac{2d_{X}}{d_{X}-1}}$ ($X=A,B$), it yields the well-known CCNR criterion [59, 13] and the entanglement criterion via SIC-POVM (ESIC) [18], respectively. For clarity and convenience, we have summarized these results in Table 1.
Table 1: Here, we summarize the relations between our criteria and these mentioned in the main text without local filtering transformation, in which CCNR and ESIC are derived from 1 with $t=\sqrt{\frac{2d}{d^{2}-1}}$ and $t=\sqrt{\frac{2d}{d-1}}$, respectively.
Criteria $M_{A}/M_{B}$ $\kappa_{A}\kappa_{B}$
Vicente [15] $\left[\begin{smallmatrix}0&\vec{0}^{\mathrm{T}}\\ \vec{0}&\mathds{1}_{d^{2}-1}\end{smallmatrix}\right]$ $\sqrt{\frac{4(d_{A}-1)(d_{B}-1)}{d_{A}d_{B}}}$
Sarbicki [19] $\left[\begin{smallmatrix}h&\vec{0}^{\mathrm{T}}\\ \vec{0}&\mathds{1}_{d^{2}-1}\end{smallmatrix}\right]$ $\sqrt{\frac{4(d_{A}-1+h_{A}^{2})(d_{B}-1+h_{B}^{2})}{d_{A}d_{B}}}$
This work $\left[\begin{smallmatrix}\frac{t}{\sqrt{2d}}&\cdots&\frac{t}{\sqrt{2d}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\sqrt{t_{A}^{2}+\frac{2d_{A}}{d_{A}+1}}\sqrt{t_{B}^{2}+\frac{2d_{B}}{d_{B}+1}}$
CCNR [59, 13] $\left[\begin{smallmatrix}\frac{1}{\sqrt{d^{2}-1}}&\cdots&\frac{1}{\sqrt{d^{2}-1}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{2d_{A}d_{B}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}$
ESIC [18] $\left[\begin{smallmatrix}\frac{1}{\sqrt{d-1}}&\cdots&\frac{1}{\sqrt{d-1}}\\ \vec{a}_{1}&\cdots&\vec{a}_{d^{2}}\end{smallmatrix}\right]$ $\frac{4d_{A}d_{B}}{\sqrt{(d_{A}^{2}-1)(d_{B}^{2}-1)}}$
To elaborate further, we compare 1 with several symbolic criteria under specific quantum states: the Horodecki’s bound entangled states with white noise $\rho_{\mathrm{H}}(s,p)$, unextendible product bases (UPB) bound entangled states with white noise $\rho_{\mathrm{UPB}}(p)$, and chessboard states $\rho_{\mathrm{CB}}$. The exact forms of these quantum states are shown in the Appendix E.